• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ramesh

    Log entry time 23:53:32 on October16,2010

    Entry number 335229

    keyword=Shift Summary

    16:00 Day shift was just over with the last kinematics with LHRS angle 34.76 deg and momentum setting 2.607 GeV/c. After that they moved the left spectrometer to a new angle 37 deg and set the central momentum to 2.406 GeV/c (No cycling of Q2 and Q3 required here since we went from a high momentum a low one). We had a problem with locking NMR. The NMR locks with hall probe B (also with hall probe E) when the regulator is off. As soon as I turn the regulator on, the hall probe goes by itself to C and the NMR gets unlocked. Alexandre forced the lock to go to probe B. After that NMR got locked.

    16:50 One of the IOCs in MCC had to be rebooted. So, it may take sometime to get beam.

    17:20 No beam yet.

    17:24 MCC tells us that we are getting beam now.
    17:34 Got beam and started taking data.

    20:52 Finished this kinematics, now going to do boiling test.

    Boiling test kinematics: lhrs: 25 deg, central p = 2.0 GeV/c

    9:30 Waiting for beam for boiling test (beam is down due to IOC problem). We have asked for harp scan of 1H04A and 1H04B harps.

    10:10 MCC was not able to produce a good harp scan with a low beam (4.5 uA). We moved the target to empty one and they did the harp scan with about 13 uA beam. Still the harp values did not meet the prescribed values (sigma x and sigma y both as 100-300 um). In the meanwhile, Brad (RC) gave me a call and told me that even if the sigmas are about 200-500 um, these are not terribely bad for the present purpose; just go with that and proceed with the run plan.

    10:14 Asked for controlled access for dvcs DAQ work.
    This will take more than two hours, and certaily the remaining time of this shift.

    Calibration runs:
    Run Type Events
    7421 Elastic LH2 36k
    7422 Elastic LH2 37k
    7423 junk (just start and stop)
    7424 Elastic LH2 37k